OpenStudy (anonymous):

Solve for x: x-6/4 =5 14 15 26 8

OpenStudy (lgbasallote):

\[\frac{x-6}{4} = 5\] multiply bot sides by 4 \[\implies x - 6 = 20\] add 6 to both sides \[\implies x - 6 + 6 = 20 + 6\] combine like terms \[\implies x = 26\] does that help?